OpenStudy (anonymous):

What is the mean of the data set? 11, 16, 15, 9, 22, 12, 13, 14, 11, 11, 19, 18 A. 11.00 B. 13.50 C. 13.75 D. 14.25

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@HaileyD

OpenStudy (anonymous):

@Adi3

OpenStudy (anonymous):

@FEARLESS_JOCEY

OpenStudy (anonymous):

@Franklin_Lapinski

OpenStudy (anonymous):

add all the number's value then divied them by total number

OpenStudy (anonymous):

how

OpenStudy (anonymous):

mean= (11+16+........)/12

OpenStudy (anonymous):

2.25

OpenStudy (anonymous):

=(11+16+ 15+ 9+ 22+ 12+ 13+ 14+ 11+ 11+ 19+ 18)/12

OpenStudy (anonymous):

14.25

OpenStudy (anonymous):

@theunkownname the answer 14.25 and @AJ01 explained how you get that

OpenStudy (anonymous):

d

OpenStudy (anonymous):

=171/12=14.25

OpenStudy (anonymous):

yes